# Local Mechanisms of Compositional Generalization in Conditional Diffusion

### v2

@@ 6 Conclusion
We proposed local conditional scores as a possible mechanism for compositional generalization. Theoretically, we proved an equivalence between conditional projective composition (a specific compositional structure) and local conditional scores (which capture both pixel- and conditional-locality); this theory extends to feature-space compositionality. Empirically, we verified that length generalization in location-conditioned CLEVR models corresponds with local conditional dependencies at high noise combined with pixel-locality at low noise. Then, we demonstrated through a causal intervention that enforcing a local architecture restores length generalization in a model that previously failed. We also offered evidence for feature-space compositionality in color-conditioned CLEVR, and preliminary evidence of compositional structure in both pixel- and feature-space in SDXL. Our results support local conditional scores as a potential mechanism of compositional generalization in conditional diffusion models, offering a lens to understand when and how generalization is achieved, and potential avenues to improve it.
@@ Acknowledgments
Appendix A Related Work
@@
Kamb & Ganguli (2024); Niedoba et al. (2024) argue that “creativity” in unconditional and class-conditional models arises from a bias toward learning local denoisers, while Lukoianov et al. (2025) challenge some of the conclusions of Kamb & Ganguli (2024) by arguing that locality arises from statistical properties of the data rather than network inductive bias. Our work builds on the idea of locality by showing that it can also be a mechanism for compositional generalization in conditional models, which none of the previous works addressed. We are agnostic about whether locality arises from data statistics or inductive bias in practice (likely both are at play); our main insight is that compositional generalization can be achieved when the network learns a local – or equivalently, compositional – structure by any means. (Our Experiment 3 suggests that local architectural interventions can promote learning the underlying the compositional structure of the data, but this structure can also be learned “naturally” as in Experiment 1.) Note also that Kamb & Ganguli (2024) propose equivariance while Niedoba et al. (2024); Lukoianov et al. (2025) omit or argue against it; in our theory and experiments equivariance is not necessary.
@@
Kadkhodaie et al. (2023) propose shrinkage in a geometry-adaptive harmonic bias as a mechanism for generalization; in the framework of our theory, this can be thought of as a bias toward sparse dependencies in a particular feature-space. Gu et al. (2023) make a experimentally study of potential causes generalization in unconditional and class-conditional settings based on the characteristics of the dataset and choices for training and model. Bertrand et al. (2025) give empirical evidence that for flow models in high-dimensions, generalization arises primarily from network inductive biases rather than noise in the flow-matching loss.
@@
Du & Kaelbling (2024); Liu et al. (2022); Bradley et al. (2025) study explicit compositions of multiple diffusion models via linear score combination, demonstrating length-generalization in CLEVR in some cases. Bradley et al. (2025) proposed Projective Composition as a definition of “correct” compositions of multiple models; here we use it to precisely characterize compositionality in a single conditional model.
@@
Despite significant evidence for the existence of disentangled feature spaces (Chen et al., 2018; Kim & Mnih, 2018; Yang et al., 2023; Locatello et al., 2019; Zhang et al., 2023; Ilharco et al., 2022) and diffusion models’ ability to learn them in some cases Karras et al. (2019); Kotovenko et al. (2019); Gatys et al. (2016); Zhu et al. (2017), feature disentanglement in diffusion representations is still not fully understood.
Layout-to-Image Diffusion Models
Our theory may help to explain the success of layout-to-image methods that use architectural locality priors (Li et al., 2023; Zheng et al., 2023; Cheng et al., 2023; 2024) or inference-time locality constraints (Dahary et al., 2024; Xue et al., 2023; Xie et al., 2023) to improve multi-object generation. Although these works primarily report improved grounding and controllability rather than explicit OOD composition, their interventions can be viewed as approximate causal tests, where increased locality leads to improved multi-object behavior, consistent with our theory and similar to the causal intervention Exp. 3 (Figure 1).
@@ Appendix B Proof of Lemma 1
(Lemma 1) Let $p$ be the CPC given by
@@
and let $s$ be the LCS given by
@@
$L_{i}^{t}(\mathcal{J})$ $=\begin{cases}\{j\}\cap\mathcal{J},\quad\text{if }i\in M_{j}\\ \emptyset,\quad\text{else}\end{cases}$
@@
To see this, we first analyze $p^{t}$ at each pixel $i$. We begin by noting that if $p$ has a pixel-space PC structure at time 0 then it has an identical PC structure at all times $t$ (Bradley et al., 2025) (because adding isotropic Gaussian noise preserves the independence between subsets).
@@ Appendix C Relaxation of Lemma 1
Claim 1 (Standard; KL divergence decreases with noise).
$\frac{\partial}{\partial t}D_{\mathrm{KL}}(N_{t}[q]||N_{t}[r])=-t\left[\left(\nabla\log\frac{N_{t}[q]}{N_{t}[r]}\right)^{2}\right]<0,\quad\text{ if }q\neq r.$
@@
(Claim 1) We want to show that
$\frac{\partial}{\partial t}D_{\mathrm{KL}}(N_{t}[q]||N_{t}[r])=-t\left[\left(\nabla\log\frac{N_{t}[q]}{N_{t}[r]}\right)^{2}\right]<0,\quad\text{ if }q\neq r.$
@@
$=-t\left[\left(\nabla\log\frac{N_{t}[q]}{N_{t}[r]}\right)^{2}\right]$
@@ Appendix D Feature-space theory
Therefore, it is not enough to learn a single transform $\mathcal{A}$ and its inverse $\mathcal{A}^{-1}$ – the network actually needs to learn a time-dependent transform/inverse pair $\mathcal{A}^{t}$, $\mathcal{A}^{-1}$ accounting for the interaction between $\mathcal{A}$ and the noising process at each time $t$, which actually depends on the (unknown) distribution $p$. How feasible this is is currently unclear.
However, there are a few special-case worth noting. First, if $\mathcal{A}$ is an orthogonal transform then we have $\mathcal{A}^{t}=\mathcal{A}$ for all $t$. Second, we can show that as the noise level increases, the non-commutativity becomes less severe:
Claim 2.
$\frac{\partial}{\partial t}D_{\mathrm{KL}}(N_{t}[\mathcal{A}\sharp p]||\mathcal{A}\sharp N_{t}[p])<0.$
Thus, if the composition structure is most important for resolving global structure at high noise levels, it may be enough to learn the single transform $\mathcal{A}$ in order to approximately exploit the compositional structure.
@@
(Claim 2) We want to show that
$\frac{\partial}{\partial t}D_{\mathrm{KL}}(N_{t}[\mathcal{A}\sharp p]||\mathcal{A}\sharp N_{t}[p])<0.$
$\frac{\partial}{\partial t}D_{\mathrm{KL}}(N_{t}[\mathcal{A}\sharp p]||\mathcal{A}\sharp N_{t}[p])$ $=\frac{\partial}{\partial t}\int N_{t}[\mathcal{A}\sharp p]\log\frac{N_{t}[\mathcal{A}\sharp p]}{\mathcal{A}\sharp N_{t}[p])}$
$=\int\frac{\partial}{\partial t}N_{t}[\mathcal{A}\sharp p]\cdot\log\frac{N_{t}[\mathcal{A}\sharp p]}{\mathcal{A}\sharp N_{t}[p])}+\int N_{t}[\mathcal{A}\sharp p]\cdot\frac{\partial}{\partial t}\log\frac{N_{t}[\mathcal{A}\sharp p]}{\mathcal{A}\sharp N_{t}[p])}$
$:=I_{1}+I_{2}$
$I_{1}$ $:=\int\frac{\partial}{\partial t}N_{t}[\mathcal{A}\sharp p]\cdot\log\frac{N_{t}[\mathcal{A}\sharp p]}{\mathcal{A}\sharp N_{t}[p])}$
$=t\int\nabla^{2}N_{t}[\mathcal{A}\sharp p]\cdot\log\frac{N_{t}[\mathcal{A}\sharp p]}{\mathcal{A}\sharp N_{t}[p])}$
$=-t\int\nabla N_{t}[\mathcal{A}\sharp p]\cdot\nabla\log\frac{N_{t}[\mathcal{A}\sharp p]}{\mathcal{A}\sharp N_{t}[p])}$
$=t\int-\frac{\nabla N_{t}[\mathcal{A}\sharp p]^{2}}{N_{t}[\mathcal{A}\sharp p]}+\frac{\nabla N_{t}[\mathcal{A}\sharp p]\nabla\mathcal{A}\sharp N_{t}[p]}{\mathcal{A}\sharp N_{t}[p]}$
$I_{2}$ $=\int N_{t}[\mathcal{A}\sharp p]\cdot\frac{\partial}{\partial t}\log\frac{N_{t}[\mathcal{A}\sharp p]}{\mathcal{A}\sharp N_{t}[p])}$
$=\int N_{t}[\mathcal{A}\sharp p]\cdot\left(\frac{\frac{\partial}{\partial t}N_{t}[\mathcal{A}\sharp p]}{N_{t}[\mathcal{A}\sharp p]}-\frac{\frac{\partial}{\partial t}\mathcal{A}\sharp N_{t}[p]}{\mathcal{A}\sharp N_{t}[p]}\right)$
$=t\int N_{t}[\mathcal{A}\sharp p]\cdot\left(\frac{\nabla^{2}N_{t}[\mathcal{A}\sharp p]}{N_{t}[\mathcal{A}\sharp p]}-\frac{\nabla^{2}\mathcal{A}\sharp N_{t}[p]}{\mathcal{A}\sharp N_{t}[p]}\right)$
$=t\int\nabla^{2}N_{t}[\mathcal{A}\sharp p]-\frac{N_{t}[\mathcal{A}\sharp p]}{\mathcal{A}\sharp N_{t}[p]}\nabla^{2}\mathcal{A}\sharp N_{t}[p]$
$=-t\int\frac{N_{t}[\mathcal{A}\sharp p]}{\mathcal{A}\sharp N_{t}[p]}\nabla^{2}\mathcal{A}\sharp N_{t}[p],\quad\text{since }\int\nabla^{2}N_{t}[\mathcal{A}\sharp p]=\nabla N_{t}[\mathcal{A}\sharp p]|_{-\infty}^{\infty}=0$
$=t\int\nabla\left(\frac{N_{t}[\mathcal{A}\sharp p]}{\mathcal{A}\sharp N_{t}[p]}\right)\nabla\mathcal{A}\sharp N_{t}[p]$
$=t\int\frac{\nabla N_{t}[\mathcal{A}\sharp p]\nabla\mathcal{A}\sharp N_{t}[p]}{\mathcal{A}\sharp N_{t}[p]}-\frac{N_{t}[\mathcal{A}\sharp p]\nabla\mathcal{A}\sharp N_{t}[p]^{2}}{\mathcal{A}\sharp N_{t}[p]^{2}}$
$\frac{\partial}{\partial t}D_{\mathrm{KL}}(N_{t}[\mathcal{A}\sharp p]||\mathcal{A}\sharp N_{t}[p])$ $=I_{1}+I_{2}$
$=-t\int\frac{\nabla N_{t}[\mathcal{A}\sharp p]^{2}}{N_{t}[\mathcal{A}\sharp p]}-2\frac{\nabla N_{t}[\mathcal{A}\sharp p]\nabla\mathcal{A}\sharp N_{t}[p]}{\mathcal{A}\sharp N_{t}[p]}+\frac{N_{t}[\mathcal{A}\sharp p]\nabla\mathcal{A}\sharp N_{t}[p]^{2}}{\mathcal{A}\sharp N_{t}[p]^{2}}$
$=-t\int N_{t}[\mathcal{A}\sharp p]\left(\frac{\nabla N_{t}[\mathcal{A}\sharp p]}{N_{t}[\mathcal{A}\sharp p]}-\frac{\nabla\mathcal{A}\sharp N_{t}[p]}{\mathcal{A}\sharp N_{t}[p]}\right)^{2}dx$
$=-t\mathbb{E}_{N_{t}[\mathcal{A}\sharp p]}\left[\left(\nabla\log\frac{N_{t}[\mathcal{A}\sharp p]}{\mathcal{A}\sharp N_{t}[p]}\right)^{2}\right]<0$
@@ Appendix E CLEVR: details and additional experiments
Table 2: Additional location-conditioning experiments. In the top table we give the maximum value, $K_{\max}$, such that the model “succeeds” for every $1\leq K\leq K_{\max}$. In the bottom table, we test $K=0$ (which is also OOD) and give the range of the number of objects typically produced. Configurations not tested are left blank. Parentheses indicate extra objects at non-conditioned locations, e.g. +(0-2) means there were 0 to 2 extra objects in addition to the $K$ at specified locations. All models are location-conditioned but with different variants: All labels means every object was labeled (Experiment 1), Single label means only one object (randomly selected) was labeled (Experiment 2), Rand num labels means a random number of objects were labeled, and Drop one label means all but one object (randomly selected) were labeled.
Train data All labels Single label Rand num labels Drop one label
1 object 1
1-2 objects 5 1 +(0-1) 3
1-3 objects 9 1 +(0-2) 5 3 +(0-1)
1-4 objects 10 1 +(0-3) 6 8 +(0-1)
1-5 objects 10 1 +(0-4) 8 9 +(0-1)
1-6 objects 11 1 +(0-5) 9 10 +(0-1)
Train data All labels Single label Rand num labels Drop one label
1 objects 0-1
1-2 objects 0 0-2 0-1
1-3 objects 0 0-3 0-3 0-1
1-4 objects 0 1-4 0-4 0-1
1-5 objects 0 0-5 0-4 0-1
1-6 objects 0 1-6 0-4 0-1

### v3

@@ 6 Conclusion
We proposed local conditional scores as a possible mechanism for compositional generalization. Theoretically, we proved an equivalence between conditional projective composition (a specific compositional structure) and local conditional scores (which capture both pixel- and conditional-locality); this theory extends to feature-space compositionality. Empirically, we verified that length generalization in location-conditioned CLEVR models corresponds with local conditional dependencies at high noise combined with pixel-locality at low noise. Then, we demonstrated through a causal intervention that enforcing a local architecture restores length generalization in a model that previously failed. We also observed length generalization in color-conditioned CLEVR, and preliminary evidence of compositional structure in both pixel- and feature-space in SDXL. Our results support local conditional scores as a potential mechanism of compositional generalization in conditional diffusion models, offering a lens to understand when and how generalization is achieved, and potential avenues to improve it.
@@ Acknowledgments
Appendix A Additional Related Work
@@
Kamb and Ganguli (2024); Niedoba et al. (2024) argue that “creativity” in unconditional and class-conditional models arises from a bias toward learning local denoisers, while Lukoianov et al. (2025) challenge some of the conclusions of Kamb and Ganguli (2024) by arguing that locality arises from statistical properties of the data rather than network inductive bias. Our work builds on the idea of locality by showing that it can also be a mechanism for compositional generalization in conditional models, which none of the previous works addressed. We are agnostic about whether locality arises from data statistics or inductive bias in practice (likely both are at play); our main insight is that compositional generalization can be achieved when the network learns a local – or equivalently, compositional – structure by any means. (Our Experiment 2L suggests that local architectural interventions can promote learning the underlying compositional structure of the data, but this structure can also be learned “naturally” as in Experiment 1.) Note also that Kamb and Ganguli (2024) propose equivariance while Niedoba et al. (2024); Lukoianov et al. (2025) omit or argue against it; in our theory and experiments equivariance is not necessary.
@@
Kadkhodaie et al. (2023) propose shrinkage in a geometry-adaptive harmonic bias as a mechanism for generalization; in the framework of our theory, this can be thought of as a bias toward sparse dependencies in a particular feature-space. Gu et al. (2023b) make an experimentally study of potential causes generalization in unconditional and class-conditional settings based on the characteristics of the dataset and choices for training and model. Bertrand et al. (2025) give empirical evidence that for flow models in high-dimensions, generalization arises primarily from network inductive biases rather than noise in the flow-matching loss.
@@
Du and Kaelbling (2024); Liu et al. (2022); Bradley et al. (2025) study explicit compositions of multiple diffusion models via linear score combination, demonstrating length-generalization in CLEVR in some cases. Bradley et al. (2025) proposed Projective Composition as a definition of “correct” compositions of multiple models; here we use it to precisely characterize compositionality in a single conditional model.
Layout-to-Image Diffusion Models
Our theory may help to explain the success of layout-to-image methods that use architectural locality priors (Li et al., 2023; Zheng et al., 2023; Cheng et al., 2023; 2024) or inference-time locality constraints (Dahary et al., 2024; Xue et al., 2023; Xie et al., 2023) to improve multi-object generation. Although these works primarily report improved grounding and controllability rather than explicit OOD composition, their interventions can be viewed as approximate causal tests, where increased locality leads to improved multi-object behavior, consistent with our theory and similar to the causal intervention Exp. 2L.
@@
There is a large body of work towards designing disentanglement metrics appropriate for “real-world” distributions (e.g. disentanglement metrics introduced by BetaVAE Higgins et al. (2017), FactorVAE Kim and Mnih (2018); MIG in Chen et al. (2018), etc.). Many works have shown evidence of disentanglement over a variety of datasets such as CelebA Karras et al. (2019); Chen et al. (2018); Kim and Mnih (2018), Shapes3D Locatello et al. (2019), and dSprites Watters et al. (2019); Chen et al. (2018); also Kotovenko et al. (2019) explores disentanglement between style and content for style transfer. Other works have shown that diffusion models have some ability to learn disentangled feature spaces Karras et al. (2019); Kotovenko et al. (2019); Gatys et al. (2016); Zhu et al. (2017). Nevertheless disentanglement in diffusion representations is still not fully understood.
@@ Appendix B Proof of Lemma 1
First, we prove that CPC $\implies$ LCS.
Let $p$ be the CPC given by
@@
Since convolution with isotropic Gaussian noise acts independently on orthogonal subspaces, this factorization is preserved at any diffusion time $t$ (Bradley et al., 2025). Thus, the time-dependent distribution is:
$p^{t}(x|c_{\mathcal{J}})$ $:=p^{t}(x_{M_{\mathcal{J}}^{\complement}}|\emptyset)\prod_{j\in\mathcal{J}}p^{t}(x_{M_{j}}|c_{j}),\quad\forall\mathcal{J}$
Let $s$ be the LCS given by
@@
$L_{i}^{t}(\mathcal{J})$ $=\begin{cases}\{j\}\cap\mathcal{J},\quad\text{if }i\in M_{j}\\ \emptyset,\quad\text{else}\end{cases}\quad N_{i}^{t}=\begin{cases}M_{j},\quad\text{if }i\in M_{j}\\ M_{b},\quad\text{else},\end{cases}\text{where }M_{b}:=M_{\mathcal{J}_{\text{all}}}^{\complement}.$
@@
To see this, we first analyze $p^{t}$ at each pixel $i$:
@@
Next, we prove the converse LCS $\implies$ CPC.
Assume $s$ is an LCS with neighborhoods $N_{i}=M_{j}$ and $L_{i}=\{j\}$ for $i\in M_{j}$. By definition, for any pixel $i\in M_{j}$, the score component $s_{i}(x|c_{\mathcal{J}})$ depends only on $x_{M_{j}}$ and $c_{j}$.
Consider the Hessian of the log-density, $H_{ik}=\frac{\partial^{2}\log p}{\partial x_{i}\partial x_{k}}=\frac{\partial s_{i}}{\partial x_{k}}$. If $i\in M_{j}$ and $k\in M_{l}$ with $j\neq l$, then $k\notin N_{i}$. By the LCS assumption, $s_{i}$ is independent of $x_{k}$, implying:
$\frac{\partial^{2}\log p}{\partial x_{i}\partial x_{k}}=0\quad\forall i\in M_{j},k\in M_{l},j\neq l.$ (3)
Since the mixed partial derivatives between disjoint sets vanish, the log-density is additively separable:
$\log p(x|c_{\mathcal{J}})=\sum_{j\in\mathcal{J}}\phi_{j}(x_{M_{j}},c_{\mathcal{J}})+C,\quad\text{with }s_{i}=\nabla_{x_{i}}\phi_{j}(x_{M_{j}},c_{\mathcal{J}})\text{ if }i\in M_{j}$ (4)
By the LCS assumption, $s_{i}$ depends only on $c_{j}$ (not the full set $c_{\mathcal{J}}$), so $\phi_{j}$ depends on conditioning only via $c_{j}$ (up to a constant). Thus:
$\log p(x|c_{\mathcal{J}})=\sum_{j\in\mathcal{J}}\tilde{\phi}_{j}(x_{M_{j}},c_{j})+\text{const}.$ (5)
Exponentiating yields $p(x|c_{\mathcal{J}})\propto\prod_{j\in\mathcal{J}}\exp(\tilde{\phi}_{j}(x_{M_{j}},c_{j}))$, which implies $p$ factorizes into independent marginals $p(x_{M_{j}}|c_{j})$. Thus, $p$ is a CPC.
@@ Appendix C Relaxation of Lemma 1
Lemma 5 (Standard; KL divergence decreases with noise).
$\frac{\partial}{\partial t}D_{\mathrm{KL}}(N_{t}[q]||N_{t}[r])=-t\mathbb{E}\left[\left(\nabla\log\frac{N_{t}[q]}{N_{t}[r]}\right)^{2}\right]<0,\quad\text{ if }q\neq r.$
@@
(Lemma 5) We want to show that
$\frac{\partial}{\partial t}D_{\mathrm{KL}}(N_{t}[q]||N_{t}[r])=-t\mathbb{E}\left[\left(\nabla\log\frac{N_{t}[q]}{N_{t}[r]}\right)^{2}\right]<0,\quad\text{ if }q\neq r.$
@@
$=-t\mathbb{E}\left[\left(\nabla\log\frac{N_{t}[q]}{N_{t}[r]}\right)^{2}\right]$
@@ Appendix D Feature-space theory
Therefore, it is not enough to learn a single transform $\mathcal{A}$ and its inverse $\mathcal{A}^{-1}$ – the network actually needs to learn a time-dependent transform/inverse pair $\mathcal{A}^{t}$, $\mathcal{A}^{-t}$ accounting for the interaction between $\mathcal{A}$ and the noising process at each time $t$, which actually depends on the (unknown) distribution $p$. How feasible this is is currently unclear.
D.1 F-LCS Heuristic
Finally, we prove Lemma 2, which provides a necessary-but-not-sufficient condition for F-LCS.
@@ D.1 F-LCS Heuristic
(Lemma 2) First, note that for scores of any distribution $p$, and any fixed choice of $x$,
[…]
$d_{i}^{t}(x)[k]$ $:=s^{t}(x|c_{i})-s^{t}(x|\emptyset)$
[…]
$\implies d_{i}^{t}(x)^{T}d_{j}^{t}(x)$ $=0,\quad\forall i\neq j,\quad\text{since }M_{i}\cap M_{j}=\emptyset,$
where in the second-to-last line we used the fact that the gradient of a function depending only on a subset of variables has zero entries in the coordinates outside that subset. The same orthogonality result also holds for $x_{0}$-parametrized networks since the score is related to the conditional mean by $\nabla\log p^{t}(x_{t}):=\frac{1}{\sigma_{t}^{2}}\mathbb{E}[x_{0}-x_{t}|x_{t}]$, therefore $d_{i}^{t}(x)\propto\mathbb{E}_{p(x_{0}|x_{t},c_{i})}[x_{0}|x_{t}]-\mathbb{E}_{p(x_{0}|x_{t},\emptyset)}[x_{0}|x_{t}]$.
Similarly, we can take an expectation over an arbitrary distribution $x\sim q$ and obtain the following orthogonality result:
[…]
$d_{i}^{t}(q)[k]$ $:=\mathbb{E}_{x\sim q}[s^{t}(x|c_{i})][k]-\mathbb{E}_{x\sim q}[s^{t}(x|\emptyset)][k]=0,\quad\forall k\notin M_{i}$
$\implies d_{i}^{t}(q)^{T}d_{j}^{t}(q)$ $=0,\quad\forall i\neq j,\quad\text{since }M_{i}\cap M_{j}=\emptyset.$
Applying the result to F-LCS scores $s_{\mathcal{A}}^{t}(z|c):=\nabla_{z}\log(\mathcal{A}\sharp p)^{t}(z|c)$, with $q\sim\mathcal{A}\sharp\mathcal{N}(0,\sigma_{t_{\max}})$, at time $t=t_{\max}$, which corresponds to evaluating the score in feature-space at the first denoising step (when the input is Gaussian noise) – gives
$d_{i}$ $:=\mathbb{E}_{\eta_{\mathcal{A}}}[s_{\mathcal{A}}^{t_{\max}}(\eta_{\mathcal{A}}|c_{i})]-\mathbb{E}_{\eta_{\mathcal{A}}}[s_{\mathcal{A}}^{t_{\max}}(\eta_{\mathcal{A}}|\emptyset)],\quad\eta_{\mathcal{A}}\sim\mathcal{A}\sharp\mathcal{N}(0,\sigma_{t_{\max}})$
$\implies d_{i}^{T}d_{j}$ $=0,\quad\forall i\neq j.$
@@ Appendix E CLEVR: details and additional experiments
Figure 7: Length-generalization in Experiments 1, 2, 3 on 1M1-M objects. We tested length-generalization from $K=0$ to $10$ conditioned locations in each model (each row shows 8 samples for a particular $K$). (Note Exp. 3 does not support $K=0$.)
Figure 8: Local causal intervention enables length generalization. (Left) Additional samples from Exp. 2L of Figure 1 (see also Section E.3). (Center and Right) Length-generalization in two different location-conditioned models, both trained on images with only a single object (and conditioned on its single location). (Center) A model with the standard EDM2 architecture does not length-generalize: it always generates exactly one object (even when conditioned on zero locations). (Right) A model with an explicitly-enforced local architecture as in Exp. 2L length-generalizes up to 5 objects, albeit with some artifacts (objects “merging” into each other). Although it does not perform as well as the Exp. 2L local model trained on 1-3 objects, any length-generalization after training on only one object is remarkable. (In Section E.2, we hypothesize that training on more objects, e.g. 1-3, may improve length-generalization by allowing the model to learn clusters of objects).
Figure 9: Additional detail for Figure 2. Locality structures in location-conditioned CLEVR models (Exps. 1, 2, 3, 2L of Figure 1). All models are conditioned on 4 locations (OOD). Each column represents a timestep $t$. Top row shows the predicted denoised images via learned scores. Lower rows (evaluated at two conditioned and one unconditioned location) show heatmaps of the pixel gradient magnitude (average absolute of Jacobian from one pixel to all other pixels), and the conditional gradient magnitude marked with $\times$ (with the “gradient” estimated via a finite difference of the score computed with and without each conditioner).
@@
Figure 11: Pixel locality structure across time for CIFAR-10, using the same EDM2 architecture from the CLEVR experiments. (Left) Each column represents a timestep $t$, and the top row shows the predicted denoised images via learned scores. Lower rows show heatmaps of the pixel gradient magnitude (average absolute Jacobian from an selected pixel to all other pixels). Evaluation at three pixel locations confirms prior empirical observations of Kamb and Ganguli (2024); Niedoba et al. (2024) that the effective local neighborhood size is large at high noise levels. (Right) Locality metrics as in Figure 12; details in Appendix F.